OpenStudy (anonymous):

Which equation(s) have the solution x = 6 (choose all that apply)? Answer options: x + 1 = 7 -x - 1 = -5 6(x + 2) = 2x - 5(x - 8) + 26 3x + 16 = 2x + 10 no explanation is necessary!

OpenStudy (anonymous):

x+1=7 x=7-1 x=6 __________________ 6(x+2)=2x-5(x-8)+26 9x=66-12 x=54/9=6

OpenStudy (anonymous):

so the answers is only x + 1 = 7? :)

OpenStudy (anonymous):

so what's the final answer @isaacme :)

OpenStudy (anonymous):

give him time! patience ;)

OpenStudy (anonymous):

realy ! Maybe calculation error, i did it faster

OpenStudy (anonymous):

ok..! ur help is appreciated though

OpenStudy (anonymous):

Wait a moment will I check

OpenStudy (anonymous):

thanks! i'll check it myself 2

OpenStudy (anonymous):

I think its correct

OpenStudy (anonymous):

yes, there are two answers .

OpenStudy (anonymous):

but isn't there supposed to be 1? @isaacme

OpenStudy (anonymous):

why? it's different equations, and in question "choose all that apply"

OpenStudy (anonymous):

yea don't mind @DaniB

OpenStudy (anonymous):

oohhhh,,, sorry didn't see that sorryyy

OpenStudy (anonymous):

its ok